# Exploring the Impact of Dataset Statistical Effect Size on Model Performance and Data Sample Size Sufficiency

### v2

@@ Results
Experiment 2 again fairly dramatically demonstrates that there is almost no correlation between effect size and data sufficiency. Experiment 2.1 suggests there is no correlation between the effect size and slope of the validation learning curve. Experiment 2.2 indicates no correlation between effect size and slope of the magnitude of the difference in error between the training and validation sets. The data points and R-Squared values suggest that determining a sufficient sample size to train your data is dependent on some other factor of our dataset along with the specific model utilized.
Discussion
Although given the limited nature of the experiments reported here, we cannot exhaustively conclude that there is no easily computable pre-hoc measure for data-sufficiency and model efficacy, it would not be surprising if this ultimately proved to be the case. Afterall, the nature of Machine Learning discards statistical certainty for predictive performance, thereby allowing for greater predictive inference than mere statistical modeling would otherwise allow. It would therefore be more surprising were such a metric easily obtainable.
Nevertheless, we believe the overriding goal of an effective pre-hoc assessment is still possible, albeit with a little bit more circuitous work needed to arrive at it.
In parallel with this work, we have conducted experiments on cross-model analysis, and have submitted for publication initial results that suggest that when comparing the internal activations of neural networks under varying conditions, models with similar behavioral architectures tend to exhibit similar levels of robustness. This suggests that if we train a model that is highly correlated to another well-established model, it will exhibit similar levels of generalizability (Oskouie, Levine, and Sarrafzadeh 2024).
This offers a potential next step in our own efforts here. If we can, from the outset, determine the likely structure of a model trained on a dataset will likely result in, and then correlate this prospective model to a preexisting one with known robustness, this may yield a useful metric on the likely robustness a model trained on a certain dataset is likely to exhibit prior to training it.
Conclusion
In this work, we attempted to demonstrate whether or not discrete statistical analyses of a dataset, in this case the Effect Size demonstrating the relative magnitude of dissonance among feature values across categorical values, could provide a useful indication of a resulting ML model’s performance or the sufficiency of the dataset size to effectively converge on an optimal model. Our results indicate that at least this particular metric bares little correlation to resulting model performance or data sufficiency. Nevertheless, we explore potential alternatives to advance the stated goal of better understanding the sufficiency of a dataset to produce a robust ML model prior to undertaking training and validation.

### v3

@@ II-B Experiment 2
Experiment 2.2 This experiment indicates no relationship between effect size and slope of the magnitude of the difference in error between the training and validation sets. The data points and metrics suggest that determining a sufficient sample size to train your data is dependent on some other factor of our dataset along with the specific model utilized. Figure 5 is consistent with the results of both experiments. On the left panel, there is a slight downtrend, though it is not prominent. The right panel reveals no observable relationship.
III Discussion
Our findings suggest that simple, univariate summaries of class separability—even when aggregated across features via mean absolute effect size—are not reliable proxies for downstream model quality or sample-size sufficiency in realistic, mixed-type tabular data. A key reason is that the “average effect size” statistic compresses a high-dimensional, heterogeneous feature set into a single scalar, discarding the structure that learning algorithms exploit. In the Adult dataset, many informative features are (i) nonlinearly expressed (e.g., interactions between education, occupation, capital gains, and hours worked) and (ii) embedded in multi-category variables with class-conditional distributions that are not well summarized by a single odds ratio. As a result, a dataset may look weakly separable when features are examined one at a time, yet still be easy to learn because of features interactions; alternatively, it may look strongly separable due to redundant or spurious correlations that do not generalize.
Moreover, the mapping from data to performance is model- and metric-dependent. Effect size is a population statistic independent of the learner, whereas generalization error depends on inductive bias, regularization, optimization, and the evaluation metric. For example, tree-based models can exploit nonlinear threshold effects and high-order feature interactions, while linear models rely more directly on marginal separability under the chosen feature representation. In this setting, a single effect-size aggregate cannot capture whether the available feature is aligned with the hypothesis class of the model family. This helps explain why correlations were small even when statistically detectable.
Taken together, our study indicates that prospective data sufficiency assessment likely requires statistics that (i) respect feature interactions and representation, and (ii) incorporate at least minimal model assumptions. Two practical directions are promising. First, replace univariate effect-size aggregation with multivariate measures of class separability that are sensitive to interactions. Second, adopt model-aware pre-hoc predictors that are still inexpensive relative to full training. These approaches preserve the goal of prospective assessment while acknowledging that “data quality” is inseparable from the learner and representation used.
More broadly, our results suggest that while simple descriptive statistics are useful for understanding marginal feature behavior, they are not reliable stand-alone predictors of model performance or sample-size sufficiency.
IV Conclusion
This paper investigated whether a simple, pre-training statistic—mean absolute feature effect size can serve as a practical heuristic for (i) predicting downstream classification performance and (ii) anticipating sample-size sufficiency via learning-curve convergence behavior on mixed-type tabular data. Across two complementary settings (varying row subsets with fixed features and varying feature sets via leave-one-feature-out) and four model families (logistic regression, decision tree, random forest, and a neural network), the observed associations between aggregated effect size and predictive metrics were consistently weak and often inconsistent. These results suggest that effect-size aggregation should not be viewed as a stand-alone predictor of model success or dataset adequacy. Future work toward data sufficiency assessment should therefore emphasize multivariate measures that preserve interaction information and low-cost, model-aware predictors that maintain practicality while improving decision relevance for experimental design and data acquisition.